OpenStudy (anonymous):

write an equation in point-slope form of the line through the point j(4,1) with slope -4

OpenStudy (anonymous):

Using point-slope form to figure this out. y1-y=m(x-x1) y-1=-4(x-4) y-1=-4x+16 +1 y=-4x=17.

OpenStudy (skullpatrol):

@halorazer the form is y-y1=m(x-x1), not y1-y=m(x-x1).

OpenStudy (skullpatrol):

$$\huge y-y_0=m(x-x_0)$$

OpenStudy (anonymous):

and that is the equation in point-slope form?...

OpenStudy (anonymous):

@skullpatrol

OpenStudy (skullpatrol):

Substitute the values given in the question into the formula.

OpenStudy (skullpatrol):

write an equation in point-slope form of the line through the point j(4,1) with slope -4 $$j(x_0, y_0)$$ slope = m

OpenStudy (skullpatrol):

What do you get @tallymark124 ?

OpenStudy (anonymous):

almost nothing on this question :(